OpenStudy (anonymous):

Please help!!! will fan and give u medals!!!!! Thank you soooooosoooooo much!!! 1.Simplify into one fraction 7/x-3+3/x-5 A.-10x-44/(x-3)(x-5) B.10x-44/(x-3)(x-4)** C.10x/(x-3)(x-5) D.10x+44/(x-3)(x-5) Simplify into one fraction 8x/x-8 - 2/x-8 A.8x+2/x-8 B.8x-2/x-8** C.3/7 D.-3/7 3.Do the data in the table represents a direct variation or an inverse variation? write an equation to model the data in the table. X-1,3,5,10 Y-4,12,20,40 A.direct variation;y=4x B.direct variation;y=1/4x C.inverse variation; xy=4 D.inverse variation; xy=1/4**

OpenStudy (anonymous):

1. \[\frac{ 2(-4x) }{ x }\]

OpenStudy (anonymous):

then \[\frac{ 10 }{ x }-8\]

OpenStudy (anonymous):

answer \[\frac{ -2(4x-5) }{ x }\]

OpenStudy (anonymous):

but that isn't an answer choice?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

It's either A or B

OpenStudy (anonymous):

what about the second one

OpenStudy (anonymous):

Yes #2 is B !

OpenStudy (anonymous):

I'm really leaning on A for #1 though

OpenStudy (anonymous):

Thank you!!! what about the last one!?

OpenStudy (anonymous):

@MelissaHolmes can you double check #1 please ?

OpenStudy (anonymous):

When two variable quantities have a constant ratio then their relationship is called a direct variation.

OpenStudy (anonymous):

so is it a or b?

OpenStudy (anonymous):

The formula for direct variation is y = kx, where k is the constant of variation.

OpenStudy (anonymous):

In an inverse variation, the values of the two variables change in an opposite way

OpenStudy (anonymous):

If the value of x is increased, then y decreases. If x decreases, the y value increases. We say that y varies inversely as the value of x.

OpenStudy (anonymous):

so its a?

OpenStudy (anonymous):

I'm trying to get you to determine which one is the answer now how did you get a ?

OpenStudy (anonymous):

idk! im so confused!